# The metric theory of small gaps for a sequence of real numbers

### v1

@@ 3.1. Proof of Theorem 1
We first arrange the elements of $A=\displaystyle\bigcup_{N\geq 1}\left(A_{N}-A_{N}\right)^{+}$ into a sequence. Let $\left(z_{n}\right)_{n\geq 1}$ be a sequence of distinct positive real numbers such that for all $N$,
@@
The sequence $\left(z_{n}\right)_{n\geq 1}$ need not be unique, since $D_{N+1}-D_{N}$ can be as large as $N$. However, the same proof works for any choice of $\left(z_{n}\right)_{n\geq 1}$ satisfying (19). Since $(a_{n})_{n\geq 1}$ is a sequence of distinct positive reals satisfying $\inf_{m\neq n}|a_{m}-a_{n}|\geq c$, we have $z_{n}\geq c$ for all $n$, and hence $\lfloor z_{n}\rfloor\geq 0$ for all $n$.
@@
Therefore, for all $\alpha\in\tilde{Q}$, and all $n\geq N_{1}(\alpha)$, we have
@@
Therefore, combining both cases (equations (31) and (32)), we obtain that for all $\alpha\in\tilde{Q}$ and all $n\geq N_{1}(\alpha)$, we have $\alpha z_{n}||\geq\psi(n)$.
@@
We now prove that for all sufficiently large $N$ and for any $\epsilon>0$, $\|\alpha z_{n^{\prime}}\|>\frac{1}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}$. We consider the finite set of indices $\{1,2,\dots,N_{1}(\alpha)-1\}$. Since the $z_{n}$ are positive and distinct, and since $\alpha$ is such that $\alpha z_{n}\notin\mathbb{Z}$ for any $n$ (which holds for almost all $\alpha$), there exists a constant $\eta(\alpha)>0$ such that
@@
Note that t= $t||=\text{ min }(\{t\},1-\{t\})\leq 1-\{$t\}, and t=t||-t||=||t|| for any $t\in\mathbb{R}$. Therefore, using (39), for all $N\geq N_{0}(\alpha)$ and all distinct $a_{m},a_{n}\in A_{N}$, we have
@@
for almost all $\alpha\in[0,1]$. The proof now follows by taking minimum over all $a_{m},a_{n}\in A_{N}$.
@@
Remark 21.
For any $\epsilon>0$, we have $\frac{\log_{2}N}{\log{N}}+(1+\epsilon)\frac{\log_{3}N}{\log{N}}=o(1)$, as $N\to\infty$. Corresponding to a fixed $\eta>0$, there exists $N_{0}^{{}^{\prime}}(\eta)$ such that $\frac{\log_{2}N+\log_{3}N^{(1+\epsilon)}}{\log{N}}<\eta$, that is, $\log N\log_{2}N^{(1+\epsilon)}<N^{\eta}$, for all $N\geq N_{0}^{{}^{\prime}}(\eta)$. Let $N_{0}(\alpha):=\max(N_{0}^{{}^{\prime}}(\alpha),N_{0}^{{}^{\prime}}(\eta))$. Since $D_{N}\leq N^{2}$, we have
for almost all $\alpha\in[0,1]$, for every $N\geq N_{0}(\alpha)$.
In other words, for all $\eta>0$, and for almost all $\alpha\in[0,1]$, we have $\delta_{\mathrm{min}}^{\alpha}(N)\geq\frac{1}{N^{2+\eta}}$ for every $N\geq N_{0}(\alpha)$. Thus, we recover [Rudnick, Theorem 1] for a sequence of distinct real numbers.
@@
Since $|f^{\prime}(x)|\geq\delta_{0}>0$ for all $x$ and $f$ is differentiable, $f$ is injective. Thus $(f(a_{n}))_{n\geq 1}$ is a sequence of distinct positive real numbers satisfying $\inf_{m\neq n}|f(a_{m})-f(a_{n})|\geq c\delta_{0}>0$. We obtain the required bounds by applying Theorem 1. ∎
@@
$\delta_{\mathrm{min}}^{\alpha}((f(n))_{n\geq 1},N)\geq\frac{1}{D_{N}\log N(\log_{2}N)^{1+\epsilon}},\text{ and}$
$\delta_{\mathrm{min}}^{\alpha}((f(n))_{n\geq 1},N)\leq 1-\frac{1}{D_{N}\log N(\log_{2}N)^{1+\epsilon}},$
@@ 3.2. Proof of Theorem 5
We first arrange the elements of $A=\displaystyle\bigcup_{N\geq 1}\left(A_{N}-A_{N}\right)^{+}$ into a sequence $\left(z_{n}\right)_{n\geq 1}$. To ensure a key monotonicity property, we choose this enumeration carefully.
For each $T\geq 2$, define the set of new differences at step $T$ by
@@
We now define the sequence $(z_{n})_{n\geq 1}$ by first listing all elements of $B_{2}$ (in any order), then all elements of $B_{3}$, and so on. This construction ensures that for all $N$, we have
@@
For each $n\in\mathbb{N}$, we define $f(n)$ to be the smallest integer $T$ such that $z_{n}\in(A_{T}-A_{T})^{+}$. By our construction, if $z_{n}\in B_{T}$, then $f(n)=T$. Since the sequence $(z_{n})_{n\geq 1}$ is ordered by the sets $B_{2},B_{3},\ldots$, the function $n\mapsto f(n)$ is non-decreasing. Also, $z_{n}\in(A_{f(n)}-A_{f(n)})^{+}$ and therefore
@@
Using (46), we obtain that the Lebesgue measure of $S_{n}$ satisfies
@@
In the next few paragraphs, we prove $\|\alpha z_{\tilde{n}}\|>\frac{c(N)}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}$ for all sufficiently large $N$. We consider the finite set of indices $\{1,2,\dots,N_{1}(\alpha)-1\}$. Since the $z_{n}$ are positive and distinct, and since $\alpha$ is such that $\alpha z_{n}$ is never an integer for any $n$ (which holds for almost all $\alpha$), there exists a constant $\eta(\alpha)>0$ such that
@@
Now, let $N_{0}(\alpha)=\max\{N_{1}(\alpha),N_{2}(\alpha)\}$. Then, for all $N\geq N_{0}(\alpha)$ and for every pair of distinct elements $a_{m},a_{n}\in A_{N}$ with corresponding index $\tilde{n}$:
@@
Note that t= $t||=\text{ min }(\{t\},1-\{t\})\leq 1-\{$t\}, and t=t||-t||=||t|| for any $t\in\mathbb{R}$. Therefore, using (60), for all $N\geq N_{0}(\alpha)$, and for all distinct $a_{m},a_{n}\in A_{N}$, we have
@@
for almost all $\alpha\in[0,1]$. The proof now follows by taking minimum over all $a_{m},a_{n}\in A_{N}$.
@@
In particular, one can take the following functions in Corollary 23 as suitable choices for $f\colon$ $\sinh(x)=\frac{e^{x}-e^{-x}}{2}$, $e^{x}$, $e^{x}+x^{k}\,(k\in\mathbb{N})$, $x^{k}\log{(x+1)}\,(k\in\mathbb{N})$, $x\log x$.
@@
Therefore, Theorem 1 applies to lacunary sequences. Since the sequence $a_{n}\coloneqq c^{n}$, where $c>1$ is a fixed constant, is lacunary, Theorem 1 also applies to such sequences-in particular, to the sequences $\left(e^{n}\right)_{n\geq 1}$, $\left(\pi^{n}\right)_{n\geq 1}$, and $\left((\frac{3}{2})^{n}\right)_{n\geq 1}$.
@@
Since $c(N)$ is a monotone decreasing sequence of positive reals, so is $d(N)$. Since $|f^{\prime}(x)|\geq\delta_{0}>0$ for all $x$ and $f$ is differentiable, $f$ is injective. Thus $(f(a_{n}))_{n\geq 1}$ is a sequence of distinct positive real numbers satisfying $\inf_{1\leq m\neq n\leq N}|f(a_{m})-f(a_{n})|\geq d(N)$. We obtain the required bound in part (a) by applying Theorem 5. Proof of part (b) follows from Theorem 5. ∎
@@
The proof follows immediately from Corollary 32 by taking $c(N)=N^{\delta-1}$, where $\delta<1$. ∎
@@ 4. Proofs of Theorems 69 and 11: Upper Bounds
The convergence part of the Borel–Cantelli lemma holds unconditionally, whereas the divergence part requires the hypothesis of pairwise independence. Consequently, proving divergence in the context of metric number theory often involves greater technical difficulty than proving convergence (see [Beresnevich-Velani] for a discussion). In Section 4.1, we briefly outline the heuristics behind the proof. In Section 4.2, we provide a proof of Theorem 6, where the upper bound depends on the size of $a_{N}$. Section 4.3.1 compiles the necessary technical lemmas and overlap estimates required for the subsequent analysis in Theorem 9, and in Section 4.3.2, we finally present the proof of Theorem 9, where the upper bound is independent of the size of $a_{N}$. In Section 4.4, we prove Theorem 11, where the upper bound result holds for all except finitely many $N$.
@@ 4.1. Outline and heuristics
Instead, we rely on the concept of “quasi-independence on average” (See Lemma 120 and cite), where the idea is to define the $S_{n}$ such that the sum of their measures diverges, while simultaneously controlling the measure of their pairwise overlaps ($S_{m}\cap S_{n}$ for $m\neq n$) to apply a suitable convergence criterion. Following Rudnick’s strategy, we employ $L^{2}$ methods that naturally lead to the study of greatest common divisor (GCD) sums, which play a central role in this area.
@@ 4.2. Proof of Theorem 6
has infinitely many solutions in $a$ and $q$. Then $\mathcal{A}$ has Lebesgue measure $1$.
@@
Let $(a_{n})_{n\geq 1}$ be a sequence of real numbers such that $\inf_{\begin{subarray}{c}1\leq m\neq n\leq N\end{subarray}}|a_{m}-a_{n}|\geq 1$ for all $N\in\mathbb{N}$. Define the finite set $A_{N}=\{a_{1},\dots,a_{N}\}$ and let $H_{N}=\#\lfloor(A_{N}-A_{N})^{+}\rfloor$. Then, $H_{N}$ satisfies the recurrence $\lceil\frac{N}{2}\rceil\leq H_{N+1}\leq H_{N}+N$. Consequently, $H_{N+1}\leq 3H_{N}+1\leq 4H_{N}$ holds for $N\geq 3$.
@@
This implies that $n\leq|a_{j}-a_{1}|,|a_{k}-a_{1}|,|a_{\ell}-a_{1}|<n+1$. By the pigeonhole principle, at least two of these points must lie on the same side of $a_{1}$ (that is, both larger or both smaller than $a_{1}$). Without loss of generality, assume both $a_{j}$ and $a_{k}$ are larger than $a_{1}$ (the case when both are smaller is similar). Then, $a_{j},a_{k}\in[a_{1}+n,a_{1}+n+1)$.
@@
Let $(a_{n})_{n\geq 1}$ be a strictly increasing sequence of real numbers such that $\inf_{\begin{subarray}{c}1\leq m\neq n\leq N\end{subarray}}|a_{m}-a_{n}|\geq 1$ for all $N\in\mathbb{N}$. Then, $H_{N}$ satisfies the recurrence $N\leq H_{N+1}\leq H_{N}+N$. Consequently, $H_{N+1}\leq 2H_{N}+1\leq 3H_{N}$ holds for $N\geq 2$.
@@
Alternatively, to show $H_{N}\geq N-1$ for an increasing sequence $(a_{n})_{n\geq 1}$ with $\inf_{m\neq n}|a_{m}-a_{n}|\geq 1$, consider the set $\mathcal{A}\coloneqq\{\lfloor a_{j}-a_{1}\rfloor:2\leq j\leq N\}$. Since $a_{j}-a_{1}=(a_{j}-a_{j-1})+(a_{j-1}-a_{1})\geq 1+(a_{j-1}-a_{1})$, we have $\lfloor a_{j}-a_{1}\rfloor\geq\lfloor a_{j-1}-a_{1}\rfloor+1$. Therefore, the values $\lfloor a_{j}-a_{1}\rfloor$ are strictly increasing with $j$, and hence all distinct. Since $\mathcal{A}\subseteq\lfloor(A_{N}-A_{N})^{+}\rfloor$, it follows that $H_{N}\geq N-1$.
@@
The sequence $(Z_{n})_{n\geq 1}$ is not uniquely defined since $H_{N+1}-H_{N}$ can be strictly larger than $1$. In fact, it can be as large as $N$, but the proof works irrespective of the version of $(Z_{n})_{n\geq 1}$ we take as long as (77) is satisfied. To reduce the size of the overlaps $S_{m}\cap S_{n}$, we use the same strategy that is used in the co-prime setup of the Duffin-Schaeffer conjecture. We replace $S_{m}$ and $S_{n}$ by modified sets that preserve most of the measure of the original sets, but remove those parts that are excessively responsible for the overlaps. To incorporate this idea into our situation, we consider the reduced sets
@@
where $\psi(n)=\frac{\log_{2}Z_{n}}{n\log n\log_{2}n}$ for all $n$. This reflects the fact that overlaps $S_{m}\cap S_{n}$ mainly come from intervals in $S_{m}$ and $S_{n}$, which are centered around points $a/Z_{m}$ and $b/Z_{n}$, respectively, for which $a/Z_{m}=$ $b/Z_{n}$, so that either $\gcd\left(a,Z_{m}\right)>1$ or $\gcd\left(b,Z_{n}\right)>1$.
@@
It is known that ${\varphi(T)/T}\gg{(\log_{2}T)^{-1}}$, that is, there exist a positive constant $C_{0}$ and an $N_{0}^{(1)}\in\mathbb{N}$ such that ${\varphi(Z_{n})/Z_{n}}\geq{C_{0}(\log_{2}Z_{n})^{-1}}$ for all $n\geq N_{0}^{(1)}$. Observe that this is a function of $Z_{n}$, meaning that it depends on the actual size of the elements of the floored positive difference set $\lfloor(A_{N}-A_{N})^{+}\rfloor$, and not a function of the cardinality of the difference set, that is, $n$.
We first assume that $\psi(n)\geq\frac{1}{2}$ for infinitely many $n$. Then, since $\|\alpha z_{n}\|\leq\frac{1}{2}$, we have $\|\alpha z_{n}\|\leq\psi(n)$ for infinitely many $n$. Otherwise, $\psi(n)\geq\frac{1}{2}$ for only finitely many $n$, that is, there exists $N_{0}^{(2)}$ such that $\psi(n)<\frac{1}{2}$ for all $n\geq N_{0}^{(2)}$, which ensures that $2\psi(n)\varphi\left(Z_{n}\right)/Z_{n}<1$. Let $N_{0}\coloneqq\max(N_{0}^{(1)},N_{0}^{(2)})$. Therefore,
@@
has infinitely many co-prime solutions $a$ and $q$.
We note that if $q\neq Z_{n}$ for some $n\in\mathbb{N}$, then $\psi^{*}(q)=0$, and inequality (84) gives $|\alpha-\frac{a}{q}|=0$, that is, $\alpha=\frac{a}{q}\in\mathbb{Q}$, which is a contradiction. In other words, $q=Z_{n}$ for some $n\in\mathbb{N}$.
@@
has infinitely many co-prime solutions $a$ and $Z_{n}$, that is, $\left\|\alpha Z_{n}\right\|\leq\frac{\log_{2}Z_{n}}{n\log n\log_{2}n}$ for infinitely many $n$. This shows that there exists a sequence $((t(n))_{n\geq 1}$ of distinct natural numbers such that $\left\|\alpha Z_{t(n)}\right\|\leq\frac{\log_{2}Z_{t(n)}}{t(n)\log t(n)\log_{2}t(n)}$ for all $n\in\mathbb{N}$.
@@
Using Lemma 38, we obtain that there are infinitely many $N$ such that
@@
We now prove that $B_{N}\leq\max(a_{1},\ldots,a_{N})\rfloor$.
@@
Since all $a_{i}>0$, we have $\min(a_{1},\ldots,a_{N})>0$, so $\lfloor|a_{m}-a_{n}|\rfloor\leq\lfloor\max(a_{1},\ldots,a_{N})\rfloor$, and we obtain the claimed upper bound for $B_{N}$.
@@
where $\lfloor\delta^{\alpha}_{\min}\rfloor(N)=\displaystyle\min_{1\leq m\neq n\leq N}\|\alpha\lfloor|a_{m}-a_{n}|\rfloor\|$, $B_{N}=\displaystyle\max_{1\leq m\neq n\leq N}\lfloor|a_{m}-a_{n}|\rfloor$ and $H_{N}=\#\lfloor(A_{N}-A_{N})^{+}\rfloor$ are defined as in Section 2.1. In particular, $B_{N}$ is upper bounded by $\lfloor a_{N}\rfloor$.
@@
We remark that although in [ABM, Theorem 1 (Equation 5)], the result is stated as
@@
(97) $\delta^{\alpha}_{\min}(N)\leq\frac{4\log_{2}\max(a_{1},\ldots,a_{N})}{D_{N}\log(\frac{N}{9})\log_{2}(\frac{N}{9})}\text{ for infinitely many }N,$
and for any increasing sequence of distinct positive integers $(a_{n})_{n\geq 1}$, the correct bound should be
@@
These follow from Corollaries 43 and 42 respectively.

### v2

@@ 3.1. Proof of Theorem 1
We first arrange the elements of $A=\displaystyle\bigcup_{N\geq 1}\left(A_{N}-A_{N}\right)^{+}$ in a sequence. Let $\left(z_{n}\right)_{n\geq 1}$ be a sequence of distinct positive real numbers such that for all $N$,
@@
The sequence $\left(z_{n}\right)_{n\geq 1}$ need not be unique, as $D_{N+1}-D_{N}$ can be as large as $N$. However, the same proof works for any choice of $\left(z_{n}\right)_{n\geq 1}$ that satisfies (19). Since $(a_{n})_{n\geq 1}$ is a sequence of distinct positive reals satisfying $\inf_{m\neq n}|a_{m}-a_{n}|\geq c$, we have $z_{n}\geq c$ for all $n$, and hence $\lfloor z_{n}\rfloor\geq 0$ for all $n$.
@@
Therefore, for all $\alpha\in\tilde{Q}$ and all $n\geq N_{1}(\alpha)$, we have
@@
Therefore, combining (31) and (32), we find that for all $\alpha\in\tilde{Q}$ and all $n\geq N_{1}(\alpha)$, we have $\alpha z_{n}||\geq\psi(n)$.
@@
We now prove that for all sufficiently large $N$ and for any $\epsilon>0$, $\|\alpha z_{n^{\prime}}\|>\frac{1}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}$. We consider a finite set of indices $\{1,2,\dots,N_{1}(\alpha)-1\}$. Since $z_{n}$ are positive and distinct, and since $\alpha$ is such that $\alpha z_{n}\notin\mathbb{Z}$ for any $n$ (which is true for almost all $\alpha$), there exists a constant $\eta(\alpha)>0$ such that
@@
Note that t= $t||=\text{ min }(\{t\},1-\{t\})\leq 1-\{$t\} and t=t||-t||=||t|| for any $t\in\mathbb{R}$. Therefore, using (39), for all $N\geq N_{0}(\alpha)$ and all distinct $a_{m},a_{n}\in A_{N}$, we have
@@
Since $|f^{\prime}(x)|\geq\delta_{0}>0$ for all $x$ and $f$ is differentiable, $f$ is injective. Thus, $(f(a_{n}))_{n\geq 1}$ is a sequence of distinct positive real numbers satisfying $\inf_{m\neq n}|f(a_{m})-f(a_{n})|\geq c\delta_{0}>0$. We obtain the required bounds by applying Theorem 1. ∎
@@
$\delta_{\mathrm{min}}^{\alpha}((f(n))_{n\geq 1},N)\geq\frac{1}{\widetilde{D}_{N}\log N(\log_{2}N)^{1+\epsilon}},\text{ and}$
$\delta_{\mathrm{min}}^{\alpha}((f(n))_{n\geq 1},N)\leq 1-\frac{1}{\widetilde{D}_{N}\log N(\log_{2}N)^{1+\epsilon}},$
@@ 3.2. Proof of Theorem 5
We first arrange the elements of $A=\displaystyle\bigcup_{N\geq 1}\left(A_{N}-A_{N}\right)^{+}$ in a sequence $\left(z_{n}\right)_{n\geq 1}$. To ensure a key monotonicity property, we choose this enumeration carefully.
For each $T\geq 2$, define the set of new differences in step $T$ by
@@
We now define the sequence $(z_{n})_{n\geq 1}$ by first listing all the elements of $B_{2}$ (in any order), then all the elements of $B_{3}$, and so on. This construction ensures that for all $N$, we have
@@
For each $n\in\mathbb{N}$, we define $f(n)$ as the smallest integer $T$ such that $z_{n}\in(A_{T}-A_{T})^{+}$. By our construction, if $z_{n}\in B_{T}$, then $f(n)=T$. Since the sequence $(z_{n})_{n\geq 1}$ is ordered by the sets $B_{2},B_{3},\ldots$, the function $n\mapsto f(n)$ is non-decreasing. Also, $z_{n}\in(A_{f(n)}-A_{f(n)})^{+}$ and therefore
@@
Using (45), we see that the Lebesgue measure of $S_{n}$ satisfies
@@
In the next few paragraphs, we prove $\|\alpha z_{\tilde{n}}\|>\frac{c(N)}{D_{N}\log N(\log_{2}N)^{1+\epsilon}}$ for all sufficiently large $N$. We consider a finite set of indices $\{1,2,\dots,N_{1}(\alpha)-1\}$. Since $z_{n}$ are positive and distinct, and since $\alpha$ is such that $\alpha z_{n}$ is never an integer for any $n$ (which holds for almost all $\alpha$), there exists a constant $\eta(\alpha)>0$ such that
@@
Now, let $N_{0}(\alpha)=\max\{N_{1}(\alpha),N_{2}(\alpha)\}$. Then, for all $N\geq N_{0}(\alpha)$ and for every pair of distinct elements $a_{m},a_{n}\in A_{N}$ with the corresponding index $\tilde{n}$:
@@
Note that t= $t||=\text{ min }(\{t\},1-\{t\})\leq 1-\{$t\} and t=t||-t||=||t|| for any $t\in\mathbb{R}$. Therefore, using (59), for all $N\geq N_{0}(\alpha)$, and for all distinct $a_{m},a_{n}\in A_{N}$, we have
@@
The following functions can be taken in Corollary 22 as suitable choices for $f(x)\colon$ $\sinh(x)=\frac{e^{x}-e^{-x}}{2}$, $e^{x}$, $e^{x}+x^{k}$, $x^{k}\log{(x+1)}$, $x\log x$, where $k\in\mathbb{N}$.
@@
Therefore, Theorem 1 applies to lacunary sequences. Since the sequence $a_{n}\coloneqq c^{n}$, where $c>1$ is a fixed constant, is lacunary, Theorem 1 also applies to such sequences, in particular, to the sequences $\left(e^{n}\right)_{n\geq 1}$, $\left(\pi^{n}\right)_{n\geq 1}$, and $\left((\frac{3}{2})^{n}\right)_{n\geq 1}$.
@@
Since $c(N)$ is a monotone decreasing sequence of positive reals, so is $d(N)$. Since $|f^{\prime}(x)|\geq\delta_{0}>0$ for all $x$ and $f$ is differentiable, $f$ is injective. Thus, $(f(a_{n}))_{n\geq 1}$ is a sequence of distinct positive real numbers satisfying $\inf_{1\leq m\neq n\leq N}|f(a_{m})-f(a_{n})|\geq d(N)$. We obtain the required bound in part (a) by applying Theorem 5. The proof of part (b) follows from Theorem 5. ∎
@@
The proof follows immediately from Corollary 31 taking $c(N)=N^{\delta-1}$, where $\delta<1$. ∎
@@ 4. Proofs of Theorems 69 and 11: Upper Bounds
The convergence part of the Borel–Cantelli lemma holds unconditionally, whereas the divergence part requires the hypothesis of pairwise independence. Consequently, proving divergence in the context of metric number theory often involves greater technical difficulty than proving convergence (see [BV2023] for a discussion). In Section 4.1, we briefly outline the heuristics behind the proof. In Section 4.2, we provide a proof of Theorem 6, where the upper bound depends on the size of $a_{N}$. Section 4.3.1 compiles the necessary technical lemmas and overlap estimates required for the subsequent analysis in Theorem 9, and in Section 4.3.2, we finally present the proof of Theorem 9, where the upper bound is independent of the size of $a_{N}$. In Section 4.4, we prove Theorem 11, where the upper bound result holds for all except finitely many $N$.
@@ 4.1. Outline and heuristics
Instead, we rely on the concept of “quasi-independence on average” (See Lemma 119 and [BDV2006, BV2010, BV2023]), where the idea is to define the $S_{n}$ such that the sum of their measures diverges, while simultaneously controlling the measure of their pairwise overlaps ($S_{m}\cap S_{n}$ for $m\neq n$) to apply a suitable convergence criterion. Following Rudnick’s strategy, we employ $L^{2}$ methods that naturally lead to the study of greatest common divisor (GCD) sums, which play a central role in this area.
@@ 4.2. Proof of Theorem 6
has infinitely many solutions in $a$ and $q$. Then $\mathcal{A}$ has the Lebesgue measure $1$.
@@
Let $(a_{n})_{n\geq 1}$ be a sequence of real numbers such that $\inf_{\begin{subarray}{c}1\leq m\neq n\leq N\end{subarray}}|a_{m}-a_{n}|\geq 1$ for all $N\in\mathbb{N}$. Define the finite set $A_{N}=\{a_{1},\dots,a_{N}\}$ and let $H_{N}=\#\lfloor(A_{N}-A_{N})^{+}\rfloor$. Then $H_{N}$ satisfies the recurrence $\lceil\frac{N}{2}\rceil\leq H_{N+1}\leq H_{N}+N$. Consequently, $H_{N+1}\leq 3H_{N}+1\leq 4H_{N}$ holds for $N\geq 3$.
@@
This implies $n\leq|a_{j}-a_{1}|,|a_{k}-a_{1}|,|a_{\ell}-a_{1}|<n+1$. By the pigeonhole principle, at least two of these points must be on the same side of $a_{1}$ (that is, both larger or both smaller than $a_{1}$). Without loss of generality, assume that both $a_{j}$ and $a_{k}$ are larger than $a_{1}$ (the case where both are smaller is similar). Then $a_{j},a_{k}\in[a_{1}+n,a_{1}+n+1)$.
@@
Let $(a_{n})_{n\geq 1}$ be a strictly increasing sequence of real numbers such that $\inf_{\begin{subarray}{c}1\leq m\neq n\leq N\end{subarray}}|a_{m}-a_{n}|\geq 1$ for all $N\in\mathbb{N}$. Then $H_{N}$ satisfies the recurrence $N\leq H_{N+1}\leq H_{N}+N$. Consequently, $H_{N+1}\leq 2H_{N}+1\leq 3H_{N}$ holds for $N\geq 2$.
@@
Alternatively, to show $H_{N}\geq N-1$ for an increasing sequence $(a_{n})_{n\geq 1}$ with $\inf_{m\neq n}|a_{m}-a_{n}|\geq 1$, consider the set $\mathcal{A}\coloneqq\{\lfloor a_{j}-a_{1}\rfloor:2\leq j\leq N\}$. Since $a_{j}-a_{1}=(a_{j}-a_{j-1})+(a_{j-1}-a_{1})\geq 1+(a_{j-1}-a_{1})$, we have $\lfloor a_{j}-a_{1}\rfloor\geq\lfloor a_{j-1}-a_{1}\rfloor+1$. Therefore, the values $\lfloor a_{j}-a_{1}\rfloor$ increase strictly with $j$, and therefore are all distinct. Since $\mathcal{A}\subseteq\lfloor(A_{N}-A_{N})^{+}\rfloor$, it follows that $H_{N}\geq N-1$.
@@
The sequence $(Z_{n})_{n\geq 1}$ is not uniquely defined, as $H_{N+1}-H_{N}$ can be strictly larger than $1$. In fact, it can be as large as $N$, but the proof works irrespective of the version of $(Z_{n})_{n\geq 1}$ we take as long as (76) is satisfied. To reduce the size of overlaps $S_{m}\cap S_{n}$, we use the same strategy as is used in the coprime setup of the Duffin-Schaeffer conjecture. We replace $S_{m}$ and $S_{n}$ with modified sets that preserve most of the measure of the original sets, but remove those parts that are excessively responsible for the overlaps. To incorporate this idea into our situation, we consider the reduced sets
@@
where $\psi(n)=\frac{\log_{2}Z_{n}}{n\log n\log_{2}n}$ for all $n$. This reflects the fact that the overlaps $S_{m}\cap S_{n}$ come mainly from intervals in $S_{m}$ and $S_{n}$, which are centred around points $a/Z_{m}$ and $b/Z_{n}$, respectively, for which $a/Z_{m}=$ $b/Z_{n}$, so that either $\gcd\left(a,Z_{m}\right)>1$ or $\gcd\left(b,Z_{n}\right)>1$.
@@
It is known that ${\varphi(T)/T}\gg{(\log_{2}T)^{-1}}$, that is, there exist a positive constant $C_{0}$ and an $N_{0}^{(1)}\in\mathbb{N}$ such that ${\varphi(Z_{n})/Z_{n}}\geq{C_{0}(\log_{2}Z_{n})^{-1}}$ for all $n\geq N_{0}^{(1)}$. Observe that this is a function of $Z_{n}$, which means that it depends on the actual size of the elements of the floored positive difference set $\lfloor(A_{N}-A_{N})^{+}\rfloor$, and not a function of the cardinality of the difference set, that is, $n$.
We first assume that $\psi(n)\geq\frac{1}{2}$ for infinitely many $n$. Since $\|\alpha z_{n}\|\leq\frac{1}{2}$, we have $\|\alpha z_{n}\|\leq\psi(n)$ for infinitely many $n$. Otherwise, $\psi(n)\geq\frac{1}{2}$ for only finitely many $n$, that is, there exists $N_{0}^{(2)}$ such that $\psi(n)<\frac{1}{2}$ for all $n\geq N_{0}^{(2)}$, which ensures that $2\psi(n)\varphi\left(Z_{n}\right)/Z_{n}<1$. Let $N_{0}\coloneqq\max(N_{0}^{(1)},N_{0}^{(2)})$. Therefore,
@@
has infinitely many coprime solutions $a$ and $q$.
We note that if $q\neq Z_{n}$ for some $n\in\mathbb{N}$, then $\psi^{*}(q)=0$, and (83) gives $|\alpha-\frac{a}{q}|=0$, that is, $\alpha=\frac{a}{q}\in\mathbb{Q}$, which is a contradiction. In other words, $q=Z_{n}$ for some $n\in\mathbb{N}$.
@@
has infinitely many coprime solutions $a$ and $Z_{n}$, that is, $\left\|\alpha Z_{n}\right\|\leq\frac{\log_{2}Z_{n}}{n\log n\log_{2}n}$ for infinitely many $n$. This shows that there exists a sequence $((t(n))_{n\geq 1}$ of distinct natural numbers such that $\left\|\alpha Z_{t(n)}\right\|\leq\frac{\log_{2}Z_{t(n)}}{t(n)\log t(n)\log_{2}t(n)}$ for all $n\in\mathbb{N}$.
@@
Using Lemma 37, we see that there are infinitely many $N$ such that
@@
We now prove that $B_{N}\leq\lfloor\max(a_{1},\ldots,a_{N})\rfloor$.
@@
Since we have all $a_{i}>0$, we have $\min(a_{1},\ldots,a_{N})>0$, so $\lfloor|a_{m}-a_{n}|\rfloor\leq\lfloor\max(a_{1},\ldots,a_{N})\rfloor$, and we obtain the upper bound claimed for $B_{N}$.
@@
where $\lfloor\delta^{\alpha}_{\min}\rfloor(N)=\displaystyle\min_{1\leq m\neq n\leq N}\|\alpha\lfloor|a_{m}-a_{n}|\rfloor\|$, $B_{N}=\displaystyle\max_{1\leq m\neq n\leq N}\lfloor|a_{m}-a_{n}|\rfloor$, and $H_{N}=\#\lfloor(A_{N}-A_{N})^{+}\rfloor$. Since $B_{N}\leq\lfloor a_{N}\rfloor$, the above statement also holds when $B_{N}$ is replaced by $\lfloor a_{N}\rfloor$.
@@
We observe that although in [ABM, Theorem 1 (Equation 5)], the result is stated as
@@
(96) $\delta^{\alpha}_{\min}(N)\leq\frac{4\log_{2}\max(a_{1},\ldots,a_{N})}{D_{N}\log(\frac{N}{9})\log_{2}(\frac{N}{9})}\text{ for infinitely many }N.$
Similarly, for any increasing sequence of distinct positive integers $(a_{n})_{n\geq 1}$, the correct bound should be
@@
These are derived from Corollaries 42 and 41, respectively.